# SpikePingpong: High-Frequency Spike Vision-based Robot Learning for Precise Striking in Table Tennis Game

### v1

@@ Abstract
Learning to control high-speed objects in the real world remains a challenging frontier in robotics. Table tennis serves as an ideal testbed for this problem, demanding both rapid interception of fast-moving balls and precise adjustment of their trajectories. This task presents two fundamental challenges: it requires a high-precision vision system capable of accurately predicting ball trajectories, and it necessitates intelligent strategic planning to ensure precise ball placement to target regions. The dynamic nature of table tennis, coupled with its real-time response requirements, makes it particularly well-suited for advancing robotic control capabilities in fast-paced, precision-critical domains. In this paper, we present SpikePingpong, a novel system that integrates spike-based vision with imitation learning for high-precision robotic table tennis. Our approach introduces two key attempts that directly address the aforementioned challenges: SONIC, a spike camera-based module that achieves millimeter-level precision in ball-racket contact prediction by compensating for real-world uncertainties such as air resistance and friction; and IMPACT, a strategic planning module that enables accurate ball placement to targeted table regions. The system harnesses a 20 kHz spike camera for high-temporal resolution ball tracking, combined with efficient neural network models for real-time trajectory correction and stroke planning. Experimental results demonstrate that SpikePingpong achieves a remarkable 91% success rate for 30 cm accuracy target area and 71% in the more challenging 20 cm accuracy task, surpassing previous state-of-the-art approaches by 38% and 37% respectively. These significant performance improvements enable the robust implementation of sophisticated tactical gameplay strategies, providing a new research perspective for robotic control in high-speed dynamic tasks.
@@ 1 Introduction
Current research in robot learning primarily focuses on manipulation tasks involving static or slow-moving objects, such as folding clothes (avigal2022speedfolding, ; wu2024unigarmentmanip, ), pouring water (wang2024gmm, ; luo2024intelligent, ), and grasping arbitrary items (vuong2024grasp, ; shao2020unigrasp, ). While these achievements represent significant progress, they address scenarios with relatively simple dynamics. Yet, the real world is replete with dynamic scenarios involving high-speed moving objects that demand rapid perception and precise control, from catching falling items (zhang2024catch, ) and intercepting projectiles (natarajan2024preprocessing, ) to navigating through crowded environments with moving obstacles (gao2023improved, ). These challenging contexts require robots to process information and react at speeds approaching or even surpassing human capabilities, thereby posing a critical frontier for advancing robotic learning beyond simplified, low-dynamic environments. To address these challenges, we explore the control of high-speed moving objects, with table tennis serving as an ideal testbed. This dynamic sport presents unique challenges beyond traditional robotics applications, requiring millisecond-level timing accuracy and millimeter-level spatial precision.
@@
When tackling high-speed dynamic tasks like robotic table tennis, existing approaches can be broadly categorized into control-based methods and learning-based methods, each with its distinct limitations in this challenging domain. Control-based approaches (acosta2003ping, ; mulling2010biomimetic, ; zhang2018real, ; silva2005robotenis, ; yang2010control, ) rely on precise physical modeling and predefined motion planning strategies. These methods benefit from mathematical rigor and computational efficiency but struggle to cope with real-world complexities, as they require precise calibration and physical parameters that make them brittle to environmental variations. Furthermore, they lack the ability to adaptively adjust strategies in response to varying ball trajectories or unexpected disturbances. Learning-based approaches (mulling2013learning, ; abeyruwan2023sim2real, ; d2023robotic, ; gao2020robotic, ; ding2022goalseye, ; zhao2024reinforcement, ), particularly reinforcement learning techniques, offer greater theoretical adaptability. However, they often suffer from the persistent sim-to-real gap, where policies trained in simulation perform poorly in physical systems. This is particularly pronounced in table tennis, where subtle factors like ball spin and contact dynamics significantly impact performance. Most similar to our work, GoalsEye (ding2022goalseye, ) employs iterative imitation learning but achieves only modest accuracy in ball placement. In contrast, our imitation learning approach circumvents the sim-to-real gap inherent in reinforcement learning while providing adaptive capabilities lacking in control-based methods, ultimately achieving significantly higher precision control than existing systems.
Figure 1: Overview of SpikePingpong. Our framework decomposes table tennis into two specialized phases: (1) the interception phase using SONIC, which leverages spike camera data to accurately intercept the incoming ball, and (2) the hitting phase employing IMPACT, which uses imitation learning to direct the ball to specified target locations. Experiments demonstrate 91% success rate for primary accuracy zones and 71% for high-precision targeting across four distinct regions.
In this paper, we propose SpikePingpong, a high-precision robotic table tennis system integrating spike-based vision and advanced control techniques as illustrated in Figure 1. Our approach decomposes the complex task of table tennis into two distinct phases with specialized technical solutions: the interception phase and the hitting phase. For interception, we introduce SONIC (Spike-Oriented Neural Improvement Calibrator), a module designed to overcome the inherent limitations of conventional control approaches. SONIC leverages a lightweight neural prediction framework trained on high-fidelity spike camera data to precisely forecast the deviation between racket center and optimal ball contact position. This innovation effectively addresses unavoidable inaccuracies in physical models caused by environmental variables and spin dynamics that are typically difficult to detect and model precisely. For hitting, our IMPACT (Imitation-based Motion Planning And Control Technology) module learns striking actions from demonstrations that map incoming ball trajectories to effective racket movements, enabling control of the ball’s landing position with remarkable precision. In summary, our contributions are as follows:
  • We propose SpikePingpong, a novel robot learning framework that decomposes the complex table tennis task into specialized interception and hitting phases, addressing the limitations of both traditional control-based and learning-based approaches.
  • We introduce a comprehensive neural control system combining SONIC, a calibration system for precise ball-racket interaction prediction, and IMPACT, an imitation learning framework that captures expert striking strategies through demonstrations to enable precise control over ball landing positions.
  • Experimental results demonstrate that SpikePingpong achieves a remarkable 91% success rate for 30cm accuracy zones and 71% for high-precision 20cm targeting across four distinct target regions, surpassing previous state-of-the-art approaches by 38% and 37% respectively.
@@ 2.1 Agile Policy Learning
Agile policy learning addresses the challenge of generating fast, adaptive, and robust behaviors in highly dynamic environments. It has been widely studied across various domains such as autonomous driving (pan2017agile, ; pomerleau1988alvinn, ; muller2005off, ; anzalone2022end, ; pan2020imitation, ), legged locomotion (nguyen2017dynamic, ; tan2018sim, ; haarnoja2018learning, ; zhong2025bridging, ), humainoid skills (he2025asap, ; ben2025homie, ; he2024omnih2o, ; zhang2024wococo, ) and dynamic manipulation tasks like throwing and catching (zhang2024catch, ; hu2023grasping, ; kim2014catching, ; huang2023dynamic, ; lan2023dexcatch, ). These tasks require policies capable of maintaining high inference frequencies, handling disturbances, and generalizing across a wide range of conditions. Perception plays a critical role in enabling robots to adapt to environmental changes (wang2022retracted, ) and to understand the dynamic interaction between objects and agents (zeng2020tossingbot, ; kober2011reinforcement, ; fu2025cordvip, ). When tracking fast-moving objects, systems often rely on high-speed motion capture setups (mori2019high, ). However, in table tennis scenarios, the high speed and abrupt motion of the ball cause significant motion blur with standard RGB cameras, leading to inaccurate position estimates and trajectory predictions. Unlike previous systems that rely on multi-camera setups, our SONIC module utilizes a neuromorphic spike camera to capture high-temporal-resolution data of ball-racket interactions. This enables us to measure and predict the deviation between expected and actual contact points, training a neural network that compensates for systematic errors without complex physical modeling. This approach significantly improves interception accuracy in our high-speed robotic table tennis system.
@@ 2.2 Robotic Table Tennis
Robotic table tennis has long served as a benchmark task in robotics due to its requirement for real-time perception, prediction, planning, and control. Ever since Billingsley initiated the first robot table tennis competition in 1983 (billingsley1983, ), the task has attracted continuous attention from the research community. The existing approaches can be broadly categorized into two groups: control-based methods and learning-based methods. Control-based approaches (acosta2003ping, ; mulling2010biomimetic, ; zhang2018real, ; silva2005robotenis, ; yang2010control, ) rely on mathematical modeling and predefined control strategies. These methods typically follow a perception-prediction-control pipeline: detecting the ball, predicting a virtual hitting point (matsushima2003learning, ; miyazaki2002realization, ), and generating joint-space trajectories to reach this point (huang2016jointly, ; koc2018online, ). While these approaches benefit from mathematical rigor, they often require precise calibration and struggle with adapting to environmental variations or unexpected ball dynamics. Learning-based methods, particularly reinforcement learning and imitation learning, have gained prominence in recent years. Reinforcement learning approaches (buchler2020learning, ; tebbe2021sample, ; gao2020robotic, ; zhao2024reinforcement, ) directly map sensory inputs to motor commands, offering greater adaptability. Abeyruwan et al. (abeyruwan2023sim2real, ) proposed an iterative sim-to-real method, while D’Aronco et al. (d2023robotic, ) presented a system addressing perception latency for transfer. Most similar to our approach, GoalsEye (ding2022goalseye, ) employs imitation learning for table tennis through demonstrations and self-supervised practice, but its reliance on sim2real transfer limits its performance. Our work takes a fundamentally different approach by employing direct imitation learning exclusively in the physical world and leveraging spike cameras for enhanced perception, achieving more precise target control capabilities without simulation limitations.
@@ 3 Method
In this section, we present our robotic table tennis framework SpikePingpong, consisting of three integrated components as shown in Figure 2. First, we employ a physics-based trajectory prediction system using a low-frequency RGB-D camera to estimate the ball’s initial hittable position. Second, we introduce SONIC (Spike-Oriented Neural Improvement Calibrator), which leverages neuromorphic vision to refine hittable position predictions by compensating for real-world physical effects. Finally, our IMPACT (Imitation-based Motion Planning And Control Technology) module generates strategic hitting motions based on visual input and landing information, enabling tactical control over return placement. This pipeline progresses from basic trajectory estimation to precise interception and strategic gameplay execution.
3.1 Ball Detection and Trajectory Prediction
@@
In high-speed robotic table tennis, we employ YOLOv4-tiny (Bubbliiiing_YOLOv4Tiny_PyTorch, ) for its computational efficiency, achieving detection frequencies up to 150 Hz. Our two-phase training strategy involves initial training on public datasets (Roboflow (alexandrova2015roboflow, ), TT2 (tt2-cyt9i_dataset, ), and Ping Pong Detection (ping-pong-detection-0guzq_dataset, )), followed by targeted fine-tuning on difficult samples with higher weights. Using an Intel RealSense D455 RGB-D camera, our system achieves over 99.8% detection accuracy, transforming 2D detections into 3D world coordinates for precise trajectory estimation and timely ball interaction.
Figure 2: Framework of SpikePingpong. Our system integrates three key components: (1) initial trajectory prediction using RGB-D camera data, (2) SONIC module for refined hittable position estimation through neuromorphic vision, and (3) IMPACT module for strategic motion planning and control. This comprehensive pipeline enables precise ball interception and tactical return placement.
Ball Trajectory Prediction.
Table tennis balls can reach velocities of 2-30m/s after being struck, requiring robotic systems to accurately predict their trajectories for successful interception. We propose a physics-based approach that enables our robotic arm to anticipate the ball’s path and position itself optimally for interception. Our prediction model first employs Exponential Moving Average filtering to obtain reliable estimates of the ball’s current position $(x,y,z)$ and velocity $(v_{x},v_{y},v_{z})$ in the world coordinate system. These filtered state estimates serve as input to our physics model, which outputs the predicted hittable position $(x_{\text{hit}},y_{\text{hit}},z_{\text{hit}})$ and corresponding velocity $(v_{x}^{\text{hit}},v_{y}^{\text{hit}},v_{z}^{\text{hit}})$. With gravitational acceleration $g$, we calculate the time $t$ required for the ball to reach the predetermined interception plane at $y_{\text{hit}}$: $t=\frac{y_{\text{hit}}-y}{v_{y}}$. Using this time value, we predict the x-coordinate at interception: $x_{\text{hit}}=x+v_{x}\cdot t$. If $x_{\text{hit}}$ falls outside the robot’s operational workspace, the ball is classified as unhittable. For the z-coordinate prediction, we consider two scenarios:
@@
While our model effectively predicts trajectories under ideal conditions—assuming constant $v_{x},v_{y}$ velocity, precise position detection, and standard gravitational acceleration $g$—real-world scenarios introduce significant deviations from these theoretical assumptions. In the following section, we describe our implementation of a high-frame-rate Spike camera system that compensates for these prediction errors, substantially enhancing the accuracy of our trajectory estimation framework.
3.2 SONIC: Spike-Oriented Neural Improvement Calibrator
Our physics-based prediction approach necessarily simplifies real-world conditions by neglecting air resistance, spin effects, and assuming constant gravitational acceleration—idealizations that introduce systematic prediction errors in practice. To address these challenges, we develop SONIC (Spike-Oriented Neural Improvement Calibrator), employ a neural network the to predict the discrepancy between the model’s predicted hittable position and the real-world hittable position.

### v2

# SpikePingpong: Spike Vision-based Fast-Slow Pingpong Robot System

@@ Abstract
Learning to control high-speed objects in dynamic environments represents a fundamental challenge in robotics. Table tennis serves as an ideal testbed for advancing robotic capabilities in dynamic environments. This task presents two fundamental challenges: it requires a high-precision vision system capable of accurately predicting ball trajectories under complex dynamics, and it necessitates intelligent control strategies to ensure precise ball striking to target regions. High-speed object manipulation typically demands advanced visual perception hardware capable of capturing rapid motion with exceptional temporal resolution. Drawing inspiration from Kahneman’s dual-system theory, where fast intuitive processing complements slower deliberate reasoning, there exists an opportunity to develop more robust perception architectures that can handle high-speed dynamics while maintaining accuracy. To this end, we present SpikePingpong, a novel system that integrates spike-based vision with imitation learning for high-precision robotic table tennis. We develop a Fast-Slow system architecture where System 1 provides rapid ball detection and preliminary trajectory prediction with millisecond-level responses, while System 2 employs spike-oriented neural calibration for precise hittable position corrections. For strategic ball striking, we introduce Imitation-based Motion Planning And Control Technology, which learns optimal robotic arm striking policies through demonstration-based learning. Experimental results demonstrate that SpikePingpong achieves a remarkable 92% success rate for 30 cm accuracy zones and 70% in the more challenging 20 cm precision targeting. This work demonstrates the potential of Fast-Slow architectures for advancing robotic capabilities in time-critical manipulation tasks.
@@ 1 Introduction
Current research in robot learning primarily focuses on manipulation tasks involving static or slow-moving objects (Avigal et al., 2022; Wu et al., 2024; Wang et al., 2024; Luo et al., 2024; Vuong et al., 2024; Shao et al., 2020). While these achievements represent significant progress, they predominantly address scenarios with relatively simple dynamics and predictable object behaviors. However, real-world environments are replete with dynamic scenarios involving high-speed moving objects that demand rapid perception and precise control, from catching falling items (Zhang et al., 2024b) and intercepting projectiles (Natarajan et al., 2024) to navigating through crowded environments (Gao et al., 2023). These high-speed scenarios present fundamentally more challenging problems requiring millisecond-level decision making and robust handling of dynamic uncertainties.
Table tennis provides an ideal testbed for developing such capabilities, as it constitutes an optimal paradigm for high-speed robotic interaction while exhibiting exceptional generalizability. This task embodies Moravec’s paradox (Moravec, 1988) in its purest form: what appears as a simple recreational activity to humans represents one of the most challenging domains for robotics, demanding the integration of high-speed perception, predictive modeling, and precise motor control. Beyond its apparent simplicity, this task systematically encapsulates the fundamental challenges of dynamic robotics: millisecond-scale perception and prediction, precise manipulation under temporal constraints, and real-time strategic planning. The core competencies developed, including high-speed object tracking, precision manipulation, and adaptive control, demonstrate direct transferability to industrial automation Deka et al. (2024), medical robotics Wah (2025), and aerospace trajectory interception systems Baradaran (2025). This inherent scalability positions table tennis as a systematic methodology for developing foundational capabilities essential for advanced robots operating in dynamic, time-critical environments.
@@
When tackling high-speed dynamic tasks like robotic table tennis, existing approaches can be broadly categorized into control-based methods and learning-based methods. Control-based approaches (Acosta et al., 2003; Mülling et al., 2010; Zhang et al., 2018; Silva et al., 2005; Yang et al., 2010) rely on precise physical modeling and predefined motion planning, but despite being mathematically rigorous and computationally efficient, they struggle with real-world complexities due to their requirement for precise calibration and inability to adaptively adjust to varying ball trajectories or unexpected disturbances. Learning-based approaches (Mülling et al., 2013; Abeyruwan et al., 2023; D’Ambrosio et al., 2023; Gao et al., 2020; Ding et al., 2022; Zhao et al., 2024; DAmbrosio et al., 2025; Su et al., 2025), offer greater theoretical adaptability. However, they often suffer from the persistent sim-to-real gap, where policies trained in simulation perform poorly in physical systems. This is particularly pronounced in table tennis, where subtle factors like ball spin and contact dynamics significantly impact performance. Additionally, existing methods typically rely on high-precision hardware vision systems, which are expensive and may still struggle with the rapid temporal dynamics required for accurate trajectory prediction. Drawing inspiration from Kahneman’s dual-system theory (Kahneman, 2011), where fast intuitive processing (System 1) complements slower deliberate reasoning (System 2), there exists an opportunity to develop more robust perception architectures that can handle high-speed dynamics while maintaining accuracy.
Figure 1: Overview of SpikePingpong. The framework integrates two key stages: (1) Interception, using a Fast-Slow architecture for precise trajectory prediction, and (2) Striking, employing the IMPACT module to execute strategic returns via imitation learning. The system achieves a 92% overall success rate and 70% in high-precision targeting tasks.
To this end, we propose SpikePingpong, a high-precision robotic table tennis system integrating spike vision-based Fast-Slow system and advanced control techniques as illustrated in Figure 1. Our system addresses the fundamental challenge of robotic table tennis through a principled decomposition into interception and striking phases, each employing specialized technical innovations. For the interception phase, we employ a Fast-Slow system architecture where System 1 provides rapid ball detection and preliminary trajectory prediction with millisecond-level responses, while System 2 leverages high-frequency spike camera data for refined trajectory analysis through neural error correction, effectively addressing physical model inaccuracies caused by environmental variables and spin dynamics. For the striking phase, we develop Imitation-based Motion Planning And Control Technology (IMPACT), which learns strategic ball striking through demonstration-based learning, mapping incoming trajectory characteristics to optimal robotic arm striking policies. In summary, our contributions are as follows:
  • We design and implement a comprehensive robotic table tennis system that systematically addresses high-speed dynamic manipulation through task-specific decomposition and Fast-Slow architecture.
  • We develop a Fast-Slow system perception framework that enables accurate trajectory prediction using conventional cameras through neural error correction, complemented by real-world imitation learning for precise ball striking control.
  • We conduct extensive experimental evaluation demonstrating superior performance with 92% success rate in 30cm zones and 70% accuracy in challenging 20cm precision targeting, validating the effectiveness of our integrated approach.
@@ 2.1 Agile Policy Learning
Agile policy learning addresses the challenge of generating fast, adaptive, and robust behaviors in highly dynamic environments. It has been widely studied across various domains such as autonomous driving (Pan et al., 2017; Pomerleau, 1988; Muller et al., 2005; Anzalone et al., 2022; Pan et al., 2020), legged locomotion (Nguyen et al., 2017; Tan et al., 2018; Haarnoja et al., 2018; Zhong et al., 2025), humanoid skills (He et al., 2025; Ben et al., 2025; He et al., 2024; Zhang et al., 2024a), and dynamic manipulation tasks like throwing and catching (Zhang et al., 2024b; Hu et al., 2023; Kim et al., 2014; Huang et al., 2023; Lan et al., 2023). These tasks require policies capable of maintaining high inference frequencies, handling disturbances, and generalizing across a wide range of conditions. Perception plays a critical role in enabling robots to adapt to environmental changes (Wang and Wang, 2022) and to understand the dynamic interaction between objects and agents (Zeng et al., 2020; Kober et al., 2011; Fu et al., 2025). When tracking fast-moving objects, systems often rely on high-speed motion capture setups (Mori et al., 2019). However, in table tennis scenarios, the high speed and abrupt motion of the ball cause significant motion blur with standard RGB cameras, leading to inaccurate position estimates and trajectory predictions. Unlike previous systems that rely on high-precision hardware setups, our SpikePingpong system introduces a Fast-Slow system architecture that integrates innovative spike-based vision technology with imitation learning. The Fast system provides rapid ball detection and preliminary trajectory estimation, while the Slow system leverages high-frequency spike camera data for refined trajectory analysis and error correction. This integrated approach compensates for systematic errors without complex physical modeling, significantly improving both interception accuracy and strategic ball striking.
@@ 2.2 Robotic Table Tennis
Robotic table tennis has long served as a benchmark task in robotics due to its requirement for real-time perception, prediction, planning, and control. Since Billingsley initiated the first robot table tennis competition in 1983 (Billingsley, 1983), the task has attracted continuous attention from the research community. Existing approaches can be broadly categorized into two groups: control-based methods and learning-based methods. Control-based approaches (Acosta et al., 2003; Mülling et al., 2010; Zhang et al., 2018) rely on mathematical modeling and predefined control strategies, typically following a perception-prediction-control pipeline. While these methods benefit from mathematical rigor, they often require precise calibration and struggle with adapting to environmental variations. Learning-based approaches, particularly reinforcement learning and imitation learning, have gained prominence recently. RL approaches (Büchler et al., 2020; Tebbe et al., 2021; Gao et al., 2020; DAmbrosio et al., 2025; Su et al., 2025) directly map sensory inputs to motor commands, offering greater adaptability. Abeyruwan et al. (Abeyruwan et al., 2023) proposed iterative sim-to-real transfer, while GoalsEye (Ding et al., 2022) employs imitation learning through demonstrations and self-supervised practice, but its sim2real reliance limits performance. Our SpikePingpong system differs by introducing a Fast-Slow system perception architecture and real-world imitation learning for ball striking, training directly on data collected from real-world interactions without complex human assistance or simulation dependencies, achieving superior performance and practical deployment.
@@ 3 Method
In this section, we present our robotic table tennis framework SpikePingpong, consisting of two integrated components as shown in Figure 2. Our approach features a Fast-Slow system for perception: System 1 employs physics-based trajectory prediction using an RGB-D camera for rapid ball detection, while System 2 leverages a high-frequency spike camera to refine predictions by compensating for real-world physical effects. For action generation, our Imitation-based Motion Planning And Control Technology (IMPACT) module generates strategic hitting motions through imitation learning, enabling tactical control over return placement.
Figure 2: Framework of SpikePingpong. The system comprises two integrated components: (1) A Fast-Slow perception architecture, where System 1 delivers rapid trajectory prediction using RGB-D data, while System 2 functions as a Spike-Oriented Neural Improvement Calibrator to refine the estimated hittable position; and (2) The IMPACT module, which facilitates strategic motion planning and control, enabling tactical return placement via imitation learning.
3.1 Fast-Slow System Ball Interception Framework
3.1.1 System 1: Rapid Ball Detection and Physics-Based Prediction.
System 1 serves as the foundation of our Fast-Slow system architecture, providing rapid response capabilities for high-speed ball tracking through two core components: real-time ball detection and physics-based trajectory prediction. The detection module extracts ball positions from RGB-D camera streams, while the prediction module utilizes classical physics models to estimate future ball trajectories and determine optimal hittable positions. Additionally, System 1 provides essential contextual information to System 2, including ball state estimates and predicted hittable position that inform the learning-based decision-making process.
@@ 3.1 Fast-Slow System Ball Interception Framework
We employ YOLOv4-tiny (bubbliiiing, 2020) for its computational efficiency, achieving detection frequencies up to 150 Hz. Our training methodology adopts a two-phase approach: initial pre-training on public datasets (Roboflow (Alexandrova et al., 2015), TT2 (desigproject, 2022), and Ping Pong Detection (pingpong, 2024)), followed by domain-specific fine-tuning. Following object detection, we perform coordinate transformation from image space to world coordinates using calibrated camera parameters, enabling accurate 3D ball position acquisition for subsequent motion planning.
Physics-Based Trajectory Prediction. Our physics-based approach represents another key component of System 1, enabling the system to anticipate the ball’s path and determine optimal hittable positions. The prediction model employs Exponential Moving Average filtering to obtain reliable estimates of the ball’s current position $(x,y,z)$ and velocity $(v_{x},v_{y},v_{z})$. These filtered state estimates serve as input to our physics model, which outputs the predicted hittable position $(x_{\text{hit}},y_{\text{hit}},z_{\text{hit}})$ and corresponding velocity $(v_{x}^{\text{hit}},v_{y}^{\text{hit}},v_{z}^{\text{hit}})$. We calculate the time $t$ required for the ball to reach the predetermined hitting plane at $y_{\text{hit}}$: $t=\frac{y_{\text{hit}}-y}{v_{y}}$. Using this time value, we predict the x-coordinate at the hittable position: $x_{\text{hit}}=x+v_{x}\cdot t$. If $x_{\text{hit}}$ falls outside the robot’s operational workspace, the ball is classified as unhittable. For the z-coordinate prediction, we consider two scenarios:
@@
3.1.2 System 2: Spike-Oriented Neural Improvement Calibrator
System 2 serves as the learning-based enhancement layer of our Fast-Slow system architecture, addressing the limitations of physics-based predictions through neural calibration. While System 1 provides rapid trajectory estimation under ideal conditions, real-world scenarios introduce deviations due to air resistance, ball spin, and sensor noise that simplified physics models cannot capture. To bridge this gap, System 2 functions as a Spike-Oriented Neural Improvement Calibrator, which learns to predict the systematic discrepancy between System 1’s theoretical hittable position and actual optimal interception position. By leveraging high-frequency spike camera observations and contextual information from System 1, System 2 provides precise calibration corrections that significantly enhance overall system accuracy.
Data Collection.